# Accelerating Edge Inference for Distributed MoE Models with Latency-Optimized Expert Placement

### v2

@@ V Related Work
Several edge serving works [li2022joint, li2023task, li2023tapfinger] have explored efficient strategies to achieve low-latency application services, yet they fail to account for the unique architectural characteristics and parallelization patterns inherent in MoE models. Recent efforts to optimize Mixture-of-Experts (MoE) models span both training and inference phases, targeting challenges such as expert placement, memory efficiency, and communication overhead. For training scenarios, FasterMoE [he2022fastermoe] optimizes training by trading off hidden-state transfers against expert migrations, while SmartMoE [zhai2023smartmoe], FlexMoE [nie2023flexmoe] and Prophet [wang2023prophet] develop GPU load balancing strategies. Lazarus [wu2024lazarus] further considers system-level scalability. However, these training-oriented approaches fail to address the heterogeneous communication and computational capabilities inherent in multi-server inference environments. For inference on memory-constrained devices, standalone systems have explored multi-level memory hierarchies and expert prefetching strategies. Methods include storing excess experts in CPU memory (as in Pre-gated MoE [hwang2024pre], which introduces pre-gating for parallel loading at potential accuracy cost), predicting expert activations to prefetch parameters (Lina [li2023accelerating], MoE-Infinity [xue2024moe], EdgeMoE [yi2023edgemoe], AdapMoE [zhong2024adapmoe], SwapMoE [kong2024swapmoe], SIDA [du2024sida]), and computing experts directly on CPU (Fiddler [kamahori2024fiddler]). While effective for individual resource-constrained devices, these solutions neglect cross-server resource coordination. In contrast, DanceMoE systematically coordinates distributed resources through optimized collaboration.
More recently, inference-time expert placement has been explored in systems like Moetuner [go2025moetuner] and DeepSeek-V3’s EPLB [deepseek2025eplb]. These methods use historical routing or expert redundancy to balance loads, but are generally designed for cluster-scale datacenter deployment, without accounting for edge-specific constraints like heterogeneous hardware or limited bandwidth. Complementary to placement strategies, communication-aware MoE systems, such as PipeMoE [shi2023pipemoe], ScheMoE [shi2024schemoe], and TUTEL [hwang2023tutel], aim to improve system efficiency through pipelining, hybrid parallelism, or fine-grained communication scheduling. While primarily designed for training, their techniques offer complementary benefits for distributed inference. In contrast to these prior efforts, DanceMoE targets collaborative inference across memory-limited edge servers. It introduces activation-aware placement and lightweight migration strategies tailored to heterogeneous devices and variable workloads—an underexplored yet increasingly important setting for scalable MoE deployment.
@@ VI Conclusion
In this paper, we present DanceMoE, a novel system designed for efficient MoE inference in edge computing environments. Unlike existing solutions tailored for cloud-scale clusters, DanceMoE is specifically architected to address the unique challenges of edge deployments, such as limited device memory, heterogeneous hardware, and inefficient network conditions. A few recent works on edge MoE propose memory-efficient strategies, targeting either a single device or homogeneous resource configurations. To this end, DanceMoE introduces adaptive expert placement and dynamic migration mechanisms that jointly optimize resource utilization and inference efficiency under workload variability. Extensive testbed experiments and simulation studies demonstrate the system’s effectiveness, adaptability to dynamic conditions, and scalability in real-world edge scenarios.

### v3

@@ V Related Work
Several edge serving works [16, 18, 17] have explored efficient strategies to achieve low-latency application services, yet they fail to account for the unique architectural characteristics and parallelization patterns inherent in MoE models. Recent efforts to optimize Mixture-of-Experts (MoE) models span both training and inference phases, targeting challenges such as expert placement, memory efficiency, and communication overhead. For training scenarios, FasterMoE [8] optimizes training by trading off hidden-state transfers against expert migrations, while SmartMoE [36], FlexMoE [23] and Prophet [29] develop GPU load balancing strategies. Lazarus [31] further considers system-level scalability. However, these training-oriented approaches fail to address the heterogeneous communication and computational capabilities inherent in multi-server inference environments. For inference on memory-constrained devices, standalone systems have explored multi-level memory hierarchies and expert prefetching strategies. Methods include storing excess experts in CPU memory (as in Pre-gated MoE [10], which introduces pre-gating for parallel loading at potential accuracy cost), predicting expert activations to prefetch parameters (Lina [14], MoE-Infinity [32], EdgeMoE [35], AdapMoE [37], SwapMoE [13], SIDA [4]), and computing experts directly on CPU (Fiddler [12]). While effective for individual resource-constrained devices, these solutions neglect cross-server resource coordination. In contrast, Prism systematically coordinates distributed resources through optimized collaboration.
More recently, inference-time expert placement has been explored in systems like Moetuner [7] and DeepSeek-V3’s EPLB [3]. These methods use historical routing or expert redundancy to balance loads, but are generally designed for cluster-scale datacenter deployment, without accounting for edge-specific constraints like heterogeneous hardware or limited bandwidth. Complementary to placement strategies, communication-aware MoE systems, such as PipeMoE [26], ScheMoE [27], and TUTEL [9], aim to improve system efficiency through pipelining, hybrid parallelism, or fine-grained communication scheduling. While primarily designed for training, their techniques offer complementary benefits for distributed inference. In contrast to these prior efforts, Prism targets collaborative inference across memory-limited edge servers. It introduces activation-aware placement and lightweight migration strategies tailored to heterogeneous devices and variable workloads—an underexplored yet increasingly important setting for scalable MoE deployment.
@@ VI Conclusion
In this paper, we present Prism, a novel system designed for efficient MoE inference in edge computing environments. Unlike existing solutions tailored for cloud-scale clusters, Prism is specifically architected to address the unique challenges of edge deployments, such as limited device memory, heterogeneous hardware, and inefficient network conditions. A few recent works on edge MoE propose memory-efficient strategies, targeting either a single device or homogeneous resource configurations. To this end, Prism introduces adaptive expert placement and dynamic migration mechanisms that jointly optimize resource utilization and inference efficiency under workload variability. Extensive testbed experiments and simulation studies demonstrate the system’s effectiveness, adaptability to dynamic conditions, and scalability in real-world edge scenarios.